# Lack of self-similarity in transverse velocity increments and circulation statistics in two-dimensional turbulence

### v1

@@ Abstract
We study the statistics of longitudinal and transverse structure functions, as well as velocity circulation in the inverse energy cascade of two-dimensional turbulence. By means of direct numerical simulations of the incompressible Navier–Stokes equations, we show that transverse structure functions exhibit an anomalous scaling, in contrast to the self-similar behavior of longitudinal ones. We argue that the intermittency of transverse structure functions explains the one observed for velocity circulation statistics, although they are driven by different turbulent structures.
Two-dimensional (2D) turbulence is a paradigmatic system in fluid dynamics as it strongly differs from three-dimensional (3D) turbulence. One of the most remarkable phenomena taking place in 2D turbulent flows is the inverse energy cascade, a process in which energy is transferred towards large scales leading to the formation of large-scale coherent structures [1, 2, 3]. This mechanism is relevant in geophysical flows such as atmospheres and oceans, which exhibit quasi-2D properties due to the suppression of motion in one direction induced by confinement, stratification, rotation, or other effects [4, 5, 6]. A prime example of this phenomenon is Jupiter’s giant red spot [7].
Unlike the direct energy cascade in three dimensions [8], it has been widely accepted that the 2D inverse energy cascade is not intermittent. More precisely, most of the attention has been devoted to the longitudinal velocity increments $\delta u^{\parallel}_{r}=(\bm{u}(\bm{x}+\bm{r})-\bm{u}(\bm{x}))\cdot\hat{r}$, which have been shown to be self-similar for a given length scale $r$ within the inertial range [9, 10]. It follows that the longitudinal structure functions (LSF) of order $p$
@@
exhibit scaling exponents $\zeta_{p}^{\parallel}$ that follow a self-similar dependence $p/3$, consistent with Kraichnan-Leith-Batchelor (KLB) theory [1, 2, 3]. Note that this is in stark contrast to 3D homogeneous and isotropic turbulence, where strong fluctuations in the velocity field lead to deviations from the Kolmogorov self-similar theory (K41) predicting $\zeta_{p}^{\parallel}=p/3$. Deviations from this scaling are typically described using multifractal theories [8].
@@
A priori, the relation between increments and circulation can be understood when writing explicitly the velocity circulation around a squared loop of opposite vertices $(x_{0},y_{0})$ and $(x_{1},y_{1})=(x_{0}+r,y_{0}+r)$
@@
We note that the integrands correspond to the transverse velocity increments in different directions. The transverse scaling exponents $\zeta_{p}^{\perp}$ are defined from the transverse structure functions (TSFs) as
@@
RUNS N $\alpha$ $\nu$ $L_{f}/L$ $\epsilon$ $\ell_{\text{I}}/L$
RUN-A 2048 $0.02$ $10^{-5}$ 256 0.001 0.135
RUN-B 6144 $0.0328$ $8\times 10^{-6}$ 768 0.033 0.272
Table 1: Parameters of the 2D numerical simulations. $N$ is the number of linear collocation points, $\alpha$ the linear friction coefficient, $\nu$ the viscosity, $L_{f}/L$ the ratio between the forcing scale and the system size, $\epsilon$ the mean energy flux of the inverse cascade, and $\ell_{\text{I}}=\int k^{-1}E(k)\mathrm{d}k/\int E(k)\mathrm{d}k$ the integral scale.
@@
where the vorticity and the stream function satisfy the relation $\omega=-\nabla^{2}\psi$, with $\nu$ the kinematic viscosity and $f$ an external forcing. The Poisson brackets are defined as $\left\{\omega,\psi\right\}=\partial_{x}\omega\partial_{y}\psi-\partial_{y}\omega\partial_{x}\psi$, and the term $-\alpha\omega$ is a linear frictional damping that dissipates energy at large scales. We analyze data from two different DNS, one provided by Guido Boffetta 2D turbulence database (RUN-A) with $N_{x}=N_{y}=N=2048$ linear collocation points [22], and a second one at a larger resolution performed by our group (RUN-B) with $N=6144$ [19]. In both cases, Eq. (5) is solved in a 2D periodic square of side $2\pi$ using a pseudo-spectral method and evolved using a second-order Adams-Bashforth (RUN-A) or Runge-Kutta (RUN-B) scheme. For RUN-A, the forcing is a random white noise in time with a Gaussian correlation in space, which satisfies $\langle f(\bm{x},t)f(0,t^{\prime})\rangle=F_{0}l_{f}^{2}\delta(t-t^{\prime})\exp[-x^{2}/(2\ell_{f}^{2})]$ with $\ell_{f}$ the forcing scale [10]. For RUN-B, we use a constant in time Gaussian forcing in Fourier space $\hat{f}(\bm{k})=\exp[-(|\bm{k}|-k_{f})^{2}/(2\Delta k^{2})]$, with $k_{f}=\ell_{f}^{-1}$ the forcing wave number and $\Delta k$ its width. We choose the parameters $\alpha$ and $k_{f}$ in order to maximize the width of the inertial range. Due to the implementation of different forcings in each simulation, the statistical properties of the 2D turbulent flow slightly differ at the forcing length scale, but become negligible in the inertial range. The turbulent properties of the flow are averaged in time once the system reaches a statistically steady state. We average 28 fields for RUN-A and 453 fields for RUN-B. The parameters of each DNS are shown in Table 1.
@@
takes negative values, signature of the energy cascading towards large scales. The superscript <k indicates that the velocity field considers only wave numbers such that $k<|\bm{k}|$ and $\epsilon=k_{f}^{-1}\sum_{k=1}^{k_{f}}\Pi(k)$ is the mean energy flux. The formation of a condensate at small wave numbers is eluded due to the damping term. Note that the energy flux is not exactly flat in the inertial range due to the use of a linear damping. It has been observed that employing an hypoviscosity term flattens the energy flux in the inertial range [22]. It is expected that in the limit of an infinite scale separation $\ell_{\alpha}/\ell_{f}\rightarrow\infty$, the flux becomes constant in a larger range of scales. Figure 1 (c) shows the energy spectra compensated by the KLB prediction $C\epsilon^{2/3}k^{-5/3}$ with $C=6$ [23].
@@
We now focus on the circulation statistics to see if there is a relation with the intermittency of TSFs. We compute the velocity circulation around squared planar loops of size $r$ over the whole space. The calculations are performed in Fourier space to preserve the spectral precision from the simulations [12]. The circulation moments $\langle|\Gamma_{r}|^{p}\rangle$ are shown in Fig. 4. The local slopes also exhibit poor scaling properties, that are improved by employing the ESS method with respect to the second-order moment $\langle|\Gamma_{r}|^{2}\rangle$. The insets of the bottom panels in Fig. 4 show the deviations of the local slopes with respect to the self-similar prediction $34p/$ 3.
@@
Despite the clear differences observed in the forcing scales between the two numerical simulations, the scaling exponents obtained from the ESS method in the inertial range are statistically equivalent. The different trend observed in the inset of Fig.5.a might be due to statistical errors and it is less than $3\%$. Remarkably, both datasets lead to the same transverse anomalous scaling. Moreover, circulation scaling exponents measured in experiments coincide well with the ones measured in both simulations without using the ESS method. These results show the robustness of the turbulent properties of the inverse cascade in 2D flows, at least when the ESS method is used.
Starting from Eq. (3) relating circulation and transverse velocity increments, it is possible to show using the Hölder inequality that (see Supplemental Material)
@@
This inequality implies that, if the TSFs are intermittent, i.e. $\zeta_{p}^{\perp}<p/3$ for $p>3$, then the circulation moments also are, as it follows that $\lambda_{p}<4p/3$ for $p>3$. This relation is obeyed in our simulations, as shown in Fig. 5.c.
We remark that in the 3D case, it has been shown that the TSFs saturate for $p\geq 10$ to a constant value [21]. Equation (8), which also holds in 3D, implies that the saturation of the $\zeta_{p}^{\perp}$ leads to $\mathrm{d}\lambda_{p}/\mathrm{d}p\leq 1$ for large $p$. This fact strongly constraints existing models of velocity circulation exponents. Assuming the bifractal fit proposed in Iyer et al. [11], the equality of Eq. (8) would take place at $p=14.5$, which is beyond the limit of statistical convergence for current numerical simulations. A similar issue arises with the multifractal model proposed in [13] at $p\approx 13$. The log-normal model proposed in [16] satisfies inequality (8) for all $p$. However, it fails for $p>33$ as the predicted exponents start to decrease.
@@ Acknowledgements.
We are grateful to Guido Boffetta for fruitful scientific discussions and for sharing his database. This work was supported by the Agence Nationale de la Recherche through the project GIANTE ANR-18-CE30-0020-01. GK was also funded by the Simons Foundation project ”Collaboration in Wave Turbulence” (award ID 651471) Computations were carried out at the Mésocentre SIGAMM hosted at the Observatoire de la Côte d’Azur.

### v2

@@ Abstract
We study the statistics of longitudinal and transverse structure functions, as well as velocity circulation, in the inverse energy cascade of two-dimensional turbulence. Using direct numerical simulations of the incompressible Navier–Stokes equations, we show that transverse structure functions exhibit an anomalous scaling, in contrast to the self-similar behavior of longitudinal ones. We derive an analytical relation that shows that the scaling exponents of transverse structure functions and velocity circulation are related in two-dimensional turbulence.
Two-dimensional (2D) turbulence is a paradigmatic system in fluid dynamics as it strongly differs from three-dimensional (3D) turbulence. One of the most remarkable phenomenon which takes place in 2D turbulent flows is the inverse energy cascade, a process in which energy is transferred towards large scales leading to the formation of large-scale coherent structures [1, 2, 3]. This mechanism is relevant in geophysical flows such as atmospheres and oceans, which exhibit quasi-2D properties due to the suppression of motion in one direction induced by confinement, stratification, rotation, or other effects [4, 5, 6]. A prime example of this phenomenon is Jupiter’s giant red spot [7].
Unlike the direct energy cascade in three dimensions [8], it has been widely accepted that the 2D inverse energy cascade is not intermittent. More precisely, most of the attention has been devoted to the longitudinal velocity increments $\delta u^{\parallel}_{r}=(\bm{u}(\bm{x}+\bm{r})-\bm{u}(\bm{x}))\cdot\hat{r}$, with $\bm{u}$ the velocity field, $\bm{x}$ the space coordinate and $\bm{r}$ the increment. This quantity has been shown to be self-similar for a given length scale $r$ within the inertial range [9, 10]. It follows that the longitudinal structure functions (LSF) of order $p$
@@
exhibit scaling exponents $\zeta_{p}^{\parallel}$ that follow a self-similar dependence $p/3$, consistent with Kraichnan-Leith-Batchelor (KLB) theory [1, 2, 3]. The angular brackets $\langle$.$\rangle$ . ⟩ indicate averaging over space and time. Note that this is in stark contrast to 3D homogeneous and isotropic turbulence, where strong fluctuations in the velocity field lead to deviations from the Kolmogorov self-similar theory (K41) predicting $\zeta_{p}^{\parallel}=p/3$. Deviations from this scaling are typically described using multifractal theories [8].
@@
A priori, the relation between increments and circulation can be understood when writing explicitly the velocity circulation around a square loop of opposite vertices $(x_{0},y_{0})$ and $(x_{1},y_{1})=(x_{0}+r,y_{0}+r)$
@@
We note that the integrands correspond to the transverse velocity increments in different directions [20]. The transverse scaling exponents $\zeta_{p}^{\perp}$ are defined from the transverse structure functions (TSFs) as
@@
RUNS N $\alpha$ $\nu$ $L/L_{f}$ $\epsilon$ $\ell_{\text{I}}/L$ $R_{\mathrm{I}}$
RUN-A 2048 $0.02$ $10^{-5}$ 256 0.001 0.135 10.61
RUN-B 6144 $0.0328$ $8\times 10^{-6}$ 768 0.033 0.272 35.2
Table 1: Parameters of the 2D numerical simulations. $N$ is the number of linear collocation points, $\alpha$ the linear friction coefficient, $\nu$ the viscosity, $L/L_{f}$ the ratio between the system size and the forcing scale, $\epsilon$ the mean energy flux of the inverse cascade, $\ell_{\text{I}}=\int k^{-1}E(k)\mathrm{d}k/\int E(k)\mathrm{d}k$ the integral scale, and $R_{\mathrm{I}}=(\ell_{\text{I}}/L_{f})^{2/3}$ a Reynolds number based on the integral and forcing scales.
@@
where the vorticity and the stream function satisfy the relation $\omega=-\nabla^{2}\psi$, with $\nu$ the kinematic viscosity and $f$ an external forcing. The Poisson brackets are defined as $\left\{\omega,\psi\right\}=\partial_{x}\omega\partial_{y}\psi-\partial_{y}\omega\partial_{x}\psi$, and the term $-\alpha\omega$ is a linear frictional damping that dissipates energy at large scales. We analyze data from two different DNS, one kindly provided by Guido Boffetta (RUN-A) with $N_{x}=N_{y}=N=2048$ linear collocation points, and a second one at a larger resolution performed by our group (RUN-B) with $N=6144$ [19]. In both cases, Eq. (5) is solved in a 2D periodic square of side $2\pi$ using a pseudo-spectral method and evolved using a second-order Adams-Bashforth (RUN-A) or Runge-Kutta (RUN-B) scheme. For RUN-A, the forcing is a random white noise in time with a Gaussian correlation in space, which satisfies $\langle f(\bm{x},t)f(0,t^{\prime})\rangle=F_{0}l_{f}^{2}\delta(t-t^{\prime})\exp[-x^{2}/(2\ell_{f}^{2})]$ with $\ell_{f}$ the forcing scale [10]. For RUN-B, we use a constant in time Gaussian forcing in Fourier space $\hat{f}(\bm{k})=\exp[-(|\bm{k}|-k_{f})^{2}/(2\Delta k^{2})]$, with $k_{f}=\ell_{f}^{-1}$ the forcing wave number and $\Delta k$ its width. We choose the parameters $\alpha$ and $k_{f}$ in order to maximize the width of the inertial range. Due to the implementation of different forcings in each simulation, the statistical properties of the 2D turbulent flow slightly differ at the forcing length scale, but become negligible in the inertial range. The turbulent properties of the flow are averaged in time once the system reaches a statistically steady state. We average 28 fields for RUN-A and 453 fields for RUN-B. The parameters of each DNS are shown in Table 1.
@@
takes negative values, signature of the energy cascading towards large scales. The superscript <k indicates that the velocity field considers only wave numbers such that $k<|\bm{k}|$ and $\epsilon=k_{f}^{-1}\sum_{k=1}^{k_{f}}\Pi(k)$ is the mean energy flux. The formation of a condensate at small wave numbers is avoided by the damping term. Note that the energy flux is not exactly flat in the inertial range due to the use of a linear damping. It has been observed that employing an hypoviscosity term flattens the energy flux in the inertial range [24]. It is expected that in the limit of an infinite scale separation $\ell_{\alpha}/\ell_{f}\rightarrow\infty$, the flux becomes constant in a larger range of scales. Figure 1 (c) shows the energy spectra compensated by the KLB prediction $C\epsilon^{2/3}k^{-5/3}$ with $C=6$ [23].
@@
We now focus on the circulation statistics to see if there is a relation with the intermittency of TSFs. We compute the velocity circulation around squared planar loops of size $r$ over the whole space. The calculations are performed in Fourier space to preserve the spectral precision from the simulations [12]. The circulation moments $\langle|\Gamma_{r}|^{p}\rangle$ are shown in Fig. 4. The local slopes also exhibit poor scaling properties, which are improved by employing the ESS method with respect to the second-order moment $\langle|\Gamma_{r}|^{2}\rangle$. The insets of the bottom panels in Fig. 4 show the deviations of the local slopes with respect to the self-similar prediction $34p/$ 3.
@@
Despite the clear differences observed in the forcing scales between the two numerical simulations, the scaling exponents obtained from the ESS method in the inertial range are statistically equivalent. The different trend observed in the inset of Fig.5a might be due to statistical errors and it is less than $3\%$. Remarkably, both datasets lead to the same transverse anomalous scaling. Moreover, circulation scaling exponents measured in experiments coincide well with the ones measured in both simulations without using the ESS method. These results show the robustness of the turbulent properties of the inverse cascade in 2D flows, at least when the ESS method is used.
Starting from Eq. (3) relating circulation and transverse velocity increments, it is possible to show using the Hölder inequality that (see Supplemental Material [20])
@@
This inequality implies that, if the TSFs are intermittent, i.e. $\zeta_{p}^{\perp}<p/3$ for $p>3$, then the circulation moments also are, as it follows that $\lambda_{p}<4p/3$ for $p>3$. This relation is obeyed in our simulations, as shown in Fig. 5.c. We remark that the derivation of Eq. (8) is valid only in the limit $r/L_{f}\gg 1$ for the inverse energy cascade. In 3D turbulence, using data extracted from Iyer et al. [11] and Iyer et al. [22], we observe that the inequality still holds, however, the demonstration of Eq. (8) needs to be modified.
@@ Acknowledgements.
We are grateful to Guido Boffetta for fruitful scientific discussions and for kindly sharing his 2D turbulence data. This work was supported by the Agence Nationale de la Recherche through the project GIANTE ANR-18-CE30-0020-01. GK was also funded by the Simons Foundation project ”Collaboration in Wave Turbulence” (award ID 651471) Computations were carried out at the Mésocentre SIGAMM hosted at the Observatoire de la Côte d’Azur.
Appendix A Appendix: Relation between circulation and transverse scaling exponents
We consider the circulation $\Gamma_{r}(\bm{s}_{0})$ around a squared loop of size $r$, with one corner of the loop placed at $\bm{s}_{0}=(x_{0},y_{0})$. It follows that
$\langle|\Gamma_{r}|^{p}\rangle=\frac{1}{V}\int|\Gamma_{r}(\bm{s}_{0})|^{p}\mathrm{d}\bm{s}_{0}\leq\frac{1}{V}\int\left[\int_{x_{0}}^{x_{0}+r}|u_{x}(x,y_{0}+r)-u_{x}(x,y_{0})|\mathrm{d}x+\int_{y_{0}}^{y_{0}+r}|u_{y}(x_{0}+r,y)-u_{y}(x_{0},y)|\mathrm{d}y\right]^{p}\mathrm{d}\bm{s}_{0},$ (9)
where we applied the triangular inequality several times. Following a similar procedure to the one in Iyer et al. [11], we now apply the Hölder inequality for each of the two integrals in the right hand side. It leads to
$\langle|\Gamma_{r}|^{p}\rangle\leq\frac{1}{V}\int\left[r^{1/q}\left(\int_{x_{0}}^{x_{0}+r}|u_{x}(x,y_{0}+r)-u_{x}(x,y_{0})|^{p}\mathrm{d}x\right)^{1/p}+r^{1/q}\left(\int_{y_{0}}^{y_{0}+r}|u_{y}(x_{0}+r,y)-u_{y}(x_{0},y)|^{p}\mathrm{d}y\right)^{1/p}\right]^{p}\mathrm{d}\bm{s}_{0}$ (10)
with $p$ and $q$ satisfying $p^{-1}+q^{-1}=1$ for $p,q>1$.
For a sufficiently large Reynolds numbers, assuming homogeneity, isotropy, and at a fixed $r$ in the inertial range, we can approximate each inner integral by $r\langle|\delta u^{\perp}_{r}|^{p}\rangle=rS_{p}^{\perp}(r)$. Fig. 6 shows the validity of this approximation in the inertial range for RUN-A. The outer integral cancels out and we obtain
$\langle|\Gamma_{r}|^{p}\rangle\leq 2^{p}r^{p/q}r^{p/p}(S_{p}^{\perp})^{p/p}=2^{p}r^{p}S_{p}^{\perp}(r).$ (11)
Finally, we use the fact that the circulation moments and TSFs follow the scaling properties $\langle|\Gamma_{r}|^{p}\rangle\sim(r/L_{f})^{\lambda_{p}}$ and $S_{p}^{\perp}\sim(r/L_{f})^{\zeta_{p}^{\perp}}$, with $L_{f}$ the forcing scale. For the inertial range of the inverse energy cascade in two-dimensional turbulence, we take the limit $r/L_{f}\gg 1$, so we obtain an inequality for the scaling exponents
$\lambda_{p}\leq\zeta_{p}^{\perp}+p.$ (12)
Figure 6: Validation of the approximation performed between Eqs.(10) and (11) for $p=2$, with $u=u_{x}$ and $v=u_{y}$. The angular brackets $\langle$.$\rangle$ . ⟩ indicate averaging in space.